### v1

# LiCoMemory: Lightweight and Cognitive Agentic Memory for Efficient Long-Term Reasoning

Abstract
Large Language Model (LLM) agents exhibit remarkable conversational and reasoning capabilities but remain constrained by limited context windows and the lack of persistent memory. Recent efforts address these limitations via external memory architectures, often employing graph-based representations, yet most adopt flat, entangled structures that intertwine semantics with topology, leading to redundant representations, unstructured retrieval, and degraded efficiency and accuracy. To resolve these issues, we propose LiCoMemory, an end-to-end agentic memory framework for real-time updating and retrieval, which introduces CogniGraph, a lightweight hierarchical graph that utilizes entities and relations as semantic indexing layers, and employs temporal and hierarchy-aware search with integrated reranking for adaptive and coherent knowledge retrieval. Experiments on long-term dialogue benchmarks, LoCoMo and LongMemEval, show that LiCoMemory not only outperforms established baselines in temporal reasoning, multi-session consistency, and retrieval efficiency, but also notably reduces update latency. Our official code and data are available at https://github.com/EverM0re/LiCoMemory.
LiCoMemory: Lightweight and Cognitive Agentic Memory for Efficient Long-Term Reasoning
Zhengjun Huang1, Zhoujin Tian1, Qintian Guo1, Fangyuan Zhang2, Yingli Zhou3, Di Jiang4, Xiaofang Zhou1 1The Hong Kong University of Science and Technology; 2Huawei Hong Kong Research Center, Hong Kong; 3The Chinese University of Hong Kong, Shenzhen; 4WeBank, Shenzhen zhuangff@cse.ust.hk, ztianaf@cse.ust.hk, qtguo@ust.hk, zhang.fangyuan@huawei.com yinglizhou@link.cuhk.edu.cn, dijiang@webank.com, zxf@cse.ust.hk
1 Introduction
Figure 1: Motivation of LiCoMemory, illustrating how LiCoMemory resolves key challenges of existing memory frameworks. For the performance graph on the bottom, radius of the circles represent the construction token consumption per dialogue.
Large Language Models (LLMs) have demonstrated remarkable advancements across a wide range of language understanding and generation tasks Achiam et al. (2023); Bai et al. (2023) and are increasingly evolving into personalized assistants with enhanced contextual reasoning capabilities Li et al. (2024). Despite their strong generalizing and reasoning abilities, LLMs remain constrained by a critical short-term memory limitation: the finite context window. Information beyond the context window cannot be effectively preserved or recalled, leading to degraded reasoning capability and reduced response accuracy in long-term conversational scenarios Hatalis et al. (2023).
To resolve these issues, early attempts to enhance the long-term memory of agents commonly adopt Retrieval-Augmented Generation (RAG) architectures, which leverage conversation history as an external knowledge source and retrieve contextually relevant information to support response generation Singh et al. (2025); Zhang et al. (2025b). Graph-based RAG further extends this paradigm by structuring conversational content into relational graphs using heuristic rules, thereby capturing semantic dependencies among historical records and improving cross-session reasoning Zhu et al. (2025). Despite their effectiveness in retrieval, these approaches often treat memory as a static component, neglecting the inherently dynamic nature of human–agent interactions, and thus lacking mechanisms for reorganization of accumulated knowledge. Moreover, such predefined retrieval and linking strategies can lead to information loss and hinder adaptive memory evolution Li et al. (2025a). Recent studies have thus shifted toward dynamic agent memory frameworks, which extend beyond traditional RAG by modeling the evolving nature of conversational data and adaptive retrieval. Mem0 Chhikara et al. (2025) initiates this shift by introducing explicit memory operations that allow agents to manage and revise stored knowledge as conversations evolve. Building on this, MIRIX Wang and Chen (2025) enhances retrieval organization through multi-granularity memory indexing and relevance fusion, improving contextual alignment and reducing redundancy in long-term reasoning. Further, Zep Rasmussen et al. represents memory as a graph to capture relational dependencies among dialogue events, promoting interpretability but at the cost of high graph-construction overhead and retrieval latency.
While these frameworks have made progress toward structured conversational memory, some key challenges still remain. (1) Coupled and redundant graph structures: Existing graph-based memory systems often intertwine semantic content with relational topology, leading to heavy, redundant, and inflexible graph representations that are difficult to adapt to dynamic human–agent interactions Chen et al. (2025). (2) Scattered and unstructured retrieval: Owing to flat architectures and unguided retrieval mechanisms, conventional memory pipelines frequently return fragmented or contextually inconsistent information, which undermines reasoning coherence and produces semantically diluted responses Xiang et al. (2025). (3) Slow update and inference: Large, monolithic graph structures incur substantial computational overhead during incremental updates and inference, limiting their applicability in real-time interaction settings. For instance, GraphRAG Edge et al. (2024a) requires up to 20 minutes for graph and community construction per dialogue and over 2 minutes of query latency Zhang et al. .
To address these challenges, we propose LiCoMemory, an end-to-end agentic memory framework that enables real-time updating and retrieval. At its core, LiCoMemory introduces CogniGraph, a lightweight and semantically aware hierarchical graph that redefines the knowledge graph as a semantic indexing layer rather than a static repository. By using graph topology as a structural scaffold instead of embedding extensive content within nodes and edges, CogniGraph indexes and organizes knowledge while linking relational structures to their original textual sources for precise and context-aware reasoning. During inference, LiCoMemory integrates a unified reranking mechanism that jointly considers semantic similarity, hierarchical structure, and temporal relevance to achieve accurate and structured retrieval. Experimental results demonstrate that LiCoMemory achieves up to 23% improvement in accuracy over the second-best baseline on established long-term dialogue benchmarks (LoCoMo Maharana et al. (2024) and LongMemEval Wu et al. (2024)), particularly on multi-session and temporal reasoning subsets, while significantly reducing input tokens and response latency, underscoring its efficiency.
Our main contributions are summarized as follows: (1) CogniGraph for semantic organization. We introduce a novel hierarchical graph structure that decouples knowledge storage from semantic organization, transforming the graph into a lightweight, update-friendly semantic index. (2) Hierarchy and temporally sensitive retrieval. LiCoMemory performs structured, top-down retrieval guided by hierarchical relations and temporal cues, ensuring coherent and contextually relevant knowledge selection. (3) Efficient and real-time memory operations. Our lightweight design enables incremental graph construction, fast updates, and low-latency inference during ongoing user–assistant interactions. Together, these components establish LiCoMemory as a unified, real-time memory system capable of retrieving higher-quality, more relevant knowledge and generating contextually grounded responses.
2 Related Work
2.1 Retrieval-Augmented Generation (RAG)
RAG has emerged as a foundational framework for augmenting LLMs with external memory Fan et al. (2024). A typical RAG pipeline first splits prior interactions or knowledge as segments, then retrieves relevant segments during inference, and provides both the user query and retrieved content to the language model for grounded generation Zhou et al. (2025). Due to the highly entangled and semantically redundant nature of conversational data, conventional RAG often retrieves overlapping or loosely related content, resulting in fragmented context and limited reasoning continuity Li et al. (2025b). To address these limitations, recent research has introduced Graph-based RAG to agent frameworks, where the historical interactions are pre-organized as a graph structure to improve retrieval and reasoning efficiency Edge et al. (2024b). Beyond flat graph representations, several frameworks further organize memory knowledge hierarchically, such as tree-based structures Zhang et al. (2025a) and community-aware architectures Wang et al. (2025). These structured organizations capture cross-session dependencies, reduce redundancy, and enhance contextual coherence, paving the way for more structured and adaptive memory systems.
2.2 Agent Memory Augmentation
While LLM-based agents demonstrate remarkable generative capabilities, they remain constrained by limited context windows and the absence of persistent memory, often resulting in inconsistent behavior across extended interactions. Agent memory augmentation therefore emerges as a promising direction to address this limitation, aiming to equip conversational agents with external memory systems that support long-term information retention, retrieval, and reasoning. LoCoMo Maharana et al. (2024) introduce a RAG-style conversational framework capable of maintaining multi-session dialogue continuity through chunk-based retrieval and coherence-aware evaluation. A-MEM Xu et al. (2025) advances the concept of agentic memory beyond passive long-term storage by introducing a self-organizing, dynamically evolving memory architecture that autonomously constructs, links, and refines knowledge representations. LongMemEval Wu et al. (2024) further advances this line of research by proposing a dedicated long-term memory model and benchmark for evaluating temporal reasoning, knowledge updating, and cross-session consistency. Mem0 Chhikara et al. (2025) employs a scalable two-phase architecture (extraction and update) that dynamically stores and retrieves salient facts using a vector database, while also offering a graph-based variant that better supports long term memory maintenance and retrieval. More recently, Zep Rasmussen et al. structures agent memory into knowledge graphs, improving retrieval relevance but suffering from high construction overhead and retrieval latency.
Figure 2: Overview of LiCoMemory workflow. Upon interaction, dialogue chunks are incrementally organized through the CogniGraph (Preliminary), a lightweight hierarchical graph linking session summaries, entity–relation triples, and dialogue chunks via cross-layer hyperlinks. New knowledge is continuously integrated and deduplicated to preserve structural consistency (Phase 1). At inference, entity extraction and hierarchical retrieval guide top-down search across graph layers (Phase 2), followed by hierarchy-temporal–semantic reranking to generate a structured prompt for retrieval-augmented generation (Phase 3).
3 Methodology
The overall workflow of LiCoMemory is shown in Figure 2. LiCoMemory initiates real-time updates and retrievals during user–assistant interactions. After each dialogue segment, the dialogue chunk with its timestamp and session ID is sent to LiCoMemory for processing, where knowledge is organized and continuously maintained through a lightweight CogniGraph optimized for incremental updates. The system either updates an existing session summary or creates a new one, extracts and deduplicates triples, and establishes cross-level links among session summaries, triples, and dialogue chunks via unique identifiers. At inference time, high-value entities are extracted from user queries in a temporally aware manner to guide top-down retrieval—from session summaries to triples and then to original chunks. Retrieved triples are re-ranked by unified session, triple, and time level relevance, and the resulting summaries, triples, and chunks are integrated into a standard prompt for augmented generation. Further workflow details are provided in the following subsections.
3.1 CogniGraph: A Lightweight and Semantically-Aware Graph Structure
Traditional graph-based memory representations often embed extensive semantic content directly within nodes and edges, resulting in entangled representations where structural topology and information content are inseparable. Such designs produce heavy and redundant graphs that hinder efficient updating, leading to unstructured retrieval outputs. To address this, we introduce CogniGraph, a lightweight and semantically aware hierarchical graph structure that redefines the role of a knowledge graph from a knowledge repository to a semantic indexing layer. Rather than functioning as a storage container for knowledge, CogniGraph employs its graph topology as a structural scaffold that organizes and indexes information across multiple granularities, thereby facilitating efficient retrieval and reasoning.
CogniGraph is composed of three interconnected layers that progressively refine the granularity of information (See Preliminay in Figure 2). 1) Session level: Each session node stores a textual summary that captures the high-level context of a user–assistant interaction. The summary also contains a set of distilled keywords (keys) that represent the central entities, topics, or temporal markers of the dialogue session. 2) Entity-relation level: This layer constitutes a lightweight knowledge graph composed of entities and relations extracted from dialogue content. Each entity node and relation edge retains only essential identifiers without verbose descriptions. Entity-relation triples are hyperlinked to their corresponding session summaries, establishing connections between fine-grained semantic relations and their contextual origins. 3) Chunk level: The lowest layer of CogniGraph stores the original dialogue chunks from which the triples were extracted. Each chunk is also hyperlinked to the triples derived from it, ensuring bidirectional traceability between raw text and structured knowledge. From top to bottom, the information granularity increases while structural abstraction decreases, forming a coherent hierarchy that aligns semantics, context, and evidence.
This hierarchical and indexing-oriented design enables CogniGraph to remain compact, easily updatable, and less redundant, while the hyperlink-based cross-layer connections ensure structured and interpretable retrieval. By organizing knowledge as a navigable semantic index rather than an overloaded repository, CogniGraph supports multi-granular reasoning, from abstract contextual understanding to fine-grained evidence retrieval, serving as the structural foundation of LiCoMemory’s retrieval module and enabling efficient, semantically grounded knowledge access for long-term conversational reasoning.
3.2 Query Processing and Integrated Rerank
To enable accurate and context-aware retrieval, LiCoMemory adopts a query processing pipeline that combines hierarchical analysis with an integrated temporally aware re-ranking mechanism (Phase 1 and 2 of Figure 2). A user query is first analyzed through entity extraction to identify salient concepts that represent the key information needs of the user. The extracted entities are then matched against the summary level of the CogniGraph. By comparing the overlap and semantic similarity between the query entities and session summary keys, LiCoMemory ranks all session summaries based on their likelihood of containing relevant information. This process yields a prioritized set of session summaries that serve as entry points for deeper retrieval within the knowledge graph. LiCoMemory then queries the entity-relation level of CogniGraph using the extracted entities as anchors to locate triples that may contain relevant contextual information. Each retrieved triple is associated with its originating session and timestamp, providing both semantic and temporal context. To compute the overall relevance of a triple, LiCoMemory integrates three complementary factors to maintain hierarchy and temporal sensitivity: session-level relevance $S_{s}$, triple-level relevance $S_{t}$ and temporal relevance. The unified semantic relevance between the query and a triple is represented by the harmonic mean of the two semantic similarities:
$S_{\text{sem}}=\frac{2S_{s}S_{t}}{S_{s}+S_{t}},$
To incorporate temporal information without overwhelming semantic relevance, we apply a Weibull-based decay function that penalizes outdated triples while retaining a long-tailed contribution for distant timestamps:
$w(\Delta\tau)=\exp\!\Big[-\Big(\frac{\Delta\tau}{\hat{\tau}}\Big)^{k}\Big],\quad 0<k<1,$
where $\Delta\tau$ denotes the time gap between the current query and the triple’s timestamp, and $\hat{\tau}$ is the median of time gaps across retrieved triples, allowing the decay to adapt to the temporal scale of each query. Finally, the overall relevance score of each triple is defined as $R(t)=S_{\text{sem}}\times w(\Delta\tau)$, which jointly captures semantic coherence and temporal recency. Because $w(\Delta\tau)\in(0,1]$, the time dimension modulates rather than dominates the semantic signal, ensuring that highly relevant but moderately older information remains retrievable. This formulation achieves a balanced trade-off between temporal adaptivity and semantic precision. The final output is a structured prompt that consolidates highly relevant session summaries, top-ranked triples, and their corresponding original dialogue chunks from which the triples were extracted. This structured representation provides the language model with a coherent and fine-grained view of the relevant knowledge, facilitating precise reasoning across different levels of granularity and enabling robust cross-session understanding.
3.3 Real-Time Interactions
Figure 3: Practical case study of LiCoMemory.
LiCoMemory supports real-time retrievals and updates throughout user-assistant interactions, which consists of two tightly coupled processes. First, the agent retrieves relevant knowledge from the existing memory graph and generates a contextually grounded response to the user query. Then, LiCoMemory performs real-time memory updates based on both the current interaction and its preceding conversational history, ensuring that the newly acquired information is seamlessly integrated into the hierarchical structure (Phase 3 of Figure 2).
Figure 3 illustrates a practical example demonstrating how LiCoMemory performs real-time retrieval and memory update during user–assistant interactions. As introduced in Section 3.2, the user query is first parsed into structured entities (e.g., Social Media Platform, Followers, Time Period), which guide hierarchical retrieval over the CogniGraph. Relevant sessions and triples are then ranked and linked back to their source dialogue chunks, providing grounded evidence for response generation. Upon the completion of this interaction, the corresponding dialogue chunk, along with its timestamp and session ID, is transmitted to LiCoMemory for incremental update processing. The system first updates the session summary with newly acquired information. For ongoing sessions, it refines associated keywords and themes to maintain temporal coherence, while for new sessions, it creates a new summary description and its keyword set to represent the contextual core of the dialogue. Following session-level processing, the system performs triple extraction to transform the current interaction into structured knowledge units. The extracted triples are then integrated into the existing entity–relation graph and hyperlinked to their source sessions. To ensure consistency and eliminate redundancy, we employ type-aware and semantic similarity matching to detect duplicate triples. For duplicates, the system links their corresponding sources as additional hyperlinks to the existing nodes instead of creating new ones, thereby maintaining a compact and coherent graph structure. Because updates and retrieval share the same CogniGraph backbone, newly added information becomes immediately available for inference without requiring a full re-indexing process. Through continuous interaction and memory refinement, the system incrementally maintain a dynamic and temporally consistent representation of user knowledge, ensuring that each response remains contextually coherent and temporally up to date.
4 Experiments
In this section we evaluate LiCoMemory on real-world datasets to assess its performance. In particular, we aim to answer the following research questions: Q1: How effective is LiCoMemory compared with existing memory paradigms? Q2: How does the components of our system affect the final result? Q3: How does LiCoMemory perform in real-time practical scenarios?
Table 1: Evaluation on long-term memory QA benchmarks utilizing different language models. The best and second-best results are highlighted. $T_{R}$ stands for query latency.
Model Method LongmemEval LoCoMo
$T_{R}$ Acc. Rec. $T_{R}$ Acc. Rec.
Llama-3.1-70B Instruct-Turbo
LoCoMo 4.51s 17.60% 22.04% 4.90s 23.63% 25.50%
Mem0 2.73s 56.80% 61.21% 3.28s 53.22% 57.05%
A-Mem 2.98s 57.40% 62.18% 3.11s 43.84% 49.17%
Zep 2.83s 60.20% 62.74% 3.50s 40.30% 51.05%
[…]
GPT-4o-mini
LoCoMo 5.34s 16.60% 21.56% 4.72s 23.87% 24.91%
Mem0 2.83s 62.60% 71.32% 2.74s 54.68% 62.31%
A-Mem 4.52s 55.00% 59.30% 4.63s 48.59% 53.82%
Zep 3.47s 58.60% 61.02% 3.51s 44.76% 46.51%
[…]
4.1 Experimental Setup
Dataset. The performance of LiCoMemory was evaluated on two public long-term memory benchmarks: LongMemEval and LoCoMo. LongMemEval Wu et al. (2024) is a comprehensive benchmark for evaluating long-term memory in conversational agents, consists of 500 questions across six types: single-session user (S.S.U.), single-session assistant (S.S.A.), single-session preference (S.S.P.), multi-session, temporal reasoning, and knowledge update. LoCoMo Maharana et al. (2024) focuses on extremely long multi-session dialogues, containing 1,986 questions in five distinct categories: single-hop, multi-hop, temporal, open-domain and adversial reasoning.
Metrics. We evaluate different methods from two aspects: response quality and efficiency. Following previous work Wu et al. (2024), we assess response quality using Accuracy (Acc.), Recall (Rec.), and F1 score. For accuracy, we adopt the same evaluation protocol in LongMemEval Wu et al. (2024), where a large language model makes binary assessments of answer correctness. Recall is defined as the proportion of ground-truth targets correctly retrieved within the top-15 items Xu et al. (2025). And F1 score is reported as the harmonic mean of accuracy and recall. Efficiency is measured by token consumption and query latency Chhikara et al. (2025), which respectively represent the total number of tokens consumed by the LLM during query processing and the time required to complete the entire response generation.
Baselines. We compare LiCoMemory with four well-established baselines. LoCoMo Maharana et al. (2024) segments entire dialogues into textual chunks and retrieves the most relevant chunks to answer questions; Zep Rasmussen et al. employs structured memory access strategies to address complex, temporally extended queries; Mem0 Chhikara et al. (2025) performs explicit in-context memory operations for interactive agent environments; A-MEM Xu et al. (2025) introduces a dynamically evolving memory architecture to link and refine knowledge representations.
Implementation Details. All methods are evaluated under the same settings. Llama-3.1-70B-Instruct-Turbo serves as the primary LLM for memory management, while BGE-M3 is adopted for text embedding to support retrieval. (citation) During answer generation, Llama-3.1-70B-Instruct-Turbo and GPT-4o-mini are utilized as generation models. The number of retrieved memory units is set to 15. All experiments are conducted on NVIDIA 80G A100 GPUs.
4.2 Main Results(RQ1)
Table 1 presents a comprehensive comparison of LiCoMemory with representative memory frameworks across two long-term dialogue benchmarks using different backbone language models. Across both Llama-3.1-70B-Instruct-Turbo and GPT-4o-mini, LiCoMemory consistently achieves the highest accuracy and recall while maintaining the lowest or near-lowest query latency, demonstrating great robustness. Specifically, on LongMemEval, it surpasses the second-best baseline (Zep) by 9.0% in accuracy and 9.7% in recall with Llama-3.1-70B, and by 11.2% and 5.3% respectively with GPT-4o-mini. Similarly, on LoCoMo, LiCoMemory outperforms Mem0 by 9.8% in accuracy and 7.5% in recall on Llama-3.1-70B, and achieves a 12.5% and 5.8% gain under GPT-4o-mini. The observed improvement in QA performance underscores the effectiveness of the proposed CogniGraph structure, while the reduced latency and retrieval volume further highlight the benefits of its precise retrieval mechanism and lightweight, efficiency-oriented graph design.
To further analyze how LiCoMemory outperforms other frameworks on long-term dialogue benchmarks, we provide a detailed breakdown of its performance across subsets with different focuses using GPT-4o-mini as the generation model, as shown in Figure 4. As illustrated in the left chart, LiCoMemory consistently surpasses the second-best baseline (Mem0) across all subsets of LoCoMo, with a large gain observed in the Temporal-Reasoning subset, where accuracy improves by 20.7%. A smaller improvement is observed on the Adversarial subset, likely due to occasional false positives arising when the correct entries are not retrieved, as reflected by the recall results in Table 1. On the LongMemEval benchmark, LiCoMemory achieves substantial gains on the Multi-Session (26.6%) and Temporal Reasoning (15.9%) subsets, highlighting the effectiveness of our CogniGraph structure and unified reranking mechanism in capturing temporal and cross-session dependencies.
Figure 4: Accuracy breakdown of LiCoMemory and baselines on subsets of LoCoMo and LongmemEval.
4.3 Ablation Study(RQ2)
Figure 5: Performance breakdown of LiCoMemory and baselines on LoCoMo and LongmemEval.
In this section, we investigate how each major component of LiCoMemory contributes to different aspects of its capability. We individually disable key modules and evaluate their effects on both Accuracy and Recall across several representative scenarios within the LongMemEval benchmark. As shown in Figure 5, removing different components results in varying degrees of performance decrease.
w/o Structured retrieval. We remove the structured retrieval mechanism of CogniGraph, performing retrieval solely over extracted triples without considering the hierarchical organization of sessions or entity–relation links. As shown in Figure 5, this simplification causes the most severe degradation among all variants, dropping from 73.8/80.6 to 51.6/56.4 on the full dataset, and especially in the multi-session setting, where performance plunges from 64.7/73.5 to 40.6/48.2. Without structured aggregation under shared contextual summaries or temporal anchors, retrieval becomes fragmented and factual grounding weakens. These results highlight that structured retrieval is essential for maintaining cross-session coherence, linking fine-grained facts to their contextual origins, and enabling robust long-term reasoning.
w/o Temporal awareness. We remove the temporal weighting mechanism in re-ranking process and perform retrieval solely based on semantic similarity, without considering timestamp-based decay and temporal information reasoning. As shown in Figure 5, in the temporal reasoning setting, accuracy and recall drop sharply from 71.4/81.3 to 48.9/53.5, and in knowledge update, performance decreases from 80.8/85.1 to 51.9/57.0. In these time-sensitive scenarios, this ablation leads to the largest performance degradation among all variants, as the model overemphasizes semantically similar but outdated triples, causing obsolete or conflicting information to be retrieved. These findings underscore the importance of temporal awareness in preserving chronological coherence and enabling reliable reasoning over dynamically evolving knowledge in long-term dialogues.
w/o Summary. We remove the summary-level weighting from ranking process, forcing the model to rely solely on fine-grained triples for relevance estimation. As shown in Figure 5, this modification consistently degrades performance across all settings, dropping from 73.8/80.6 to 61.4/69.5 on the full dataset and from 64.7/73.5 to 42.9/51.3 in the multi-session case. Without the guidance of high-level summaries, retrieval becomes overly local and miss broader contextual cues that link related sessions and entities. These observations indicate that summary plays a crucial role in abstracting long-term context and steering the retrieval process toward semantically coherent evidence.
4.4 Detailed Case Study(RQ3)
Following the case discussed in Section 3.3, we now conduct a detailed evaluation of how LiCoMemory performs in a practical case, where a user engages with the agent and demands real-time update and retrieval from LiCoMemory. For this experiment, we use Llama-3.1-70B-Instruct-Turbo as the backbone model and perform a chunk by chunk insertion on the LongmemEval dataset to simulate the real life cases. During the insertion of the context dialogues, we do not request retrieval from CogniGraph but instead only insert the dialogue chunks, and perform retrieval for the query. We collect the data and present it as follows. $K_{G}$ denotes the average token consumption per session during the context processing stage, reflecting the system’s efficiency in summarizing and structuring dialogue history. $K_{R}$ represents the number of input tokens in the retrieval prompt, indicating the compactness of the information fed into the language model during response generation. $T_{G}$ measures the latency per session for context processing, while $T_{R}$ quantifies the query latency at inference time.As shown, LiCoMemory achieves an accuracy of 67.4%, with only 1.8% of accuracy drop compared to the static scenario, demonstrating strong robustness. Meanwhile LiCoMemory computes with modest token usage (13.52k for $K_{G}$ and 1.22k for $K_{R}$) and low latency (5.21 s for $T_{G}$ and 2.63 s for $T_{R}$), demonstrating its capability to maintain both retrieval accuracy and real-time responsiveness in dynamic conversational scenarios.
5 Conclusion
This paper presents LiCoMemory, an end-to-end agentic memory framework designed for real-time updating, retrieval, and reasoning in long-term conversational scenarios. LiCoMemory incorporates CogniGraph, a lightweight and semantically aware hierarchical graph structure that redefines the role of knowledge graphs as a semantic indexing layer rather than a static repository. By leveraging hierarchical and temporally sensitive retrieval, the system jointly considers session-level, relational-level, and temporal relevance to retrieve coherent and contextually aligned knowledge. Experimental results on long-term dialogue benchmarks demonstrate that LiCoMemory consistently retrieves high-quality information and achieves superior performance in temporal reasoning, multi-session reasoning and other complicated tasks, while significantly improving update efficiency and inference speed compared to existing memory baselines. In future work, we plan to extend our structure to multi-agent settings and explore adaptive memory compression strategies to further enhance scalability and reasoning capability.